# Continual Action Assessment via Task-Consistent Score-Discriminative Feature Distribution Modeling

### v4

@@ Abstract
Action Quality Assessment (AQA) is a task that tries to answer how well an action is carried out. While remarkable progress has been achieved, existing works on AQA assume that all the training data are visible for training in one time, but do not enable continual learning on assessing new technical actions. In this work, we address such a Continual Learning problem in AQA (Continual-AQA), which urges a unified model to learn AQA tasks sequentially without forgetting. Our idea for modeling Continual-AQA is to sequentially learn a task-consistent score-discriminative feature distribution, in which the latent features express a strong correlation with the score labels regardless of the task or action types. From this perspective, we aim to mitigate the forgetting in Continual-AQA from two aspects. Firstly, to fuse the features of new and previous data into a score-discriminative distribution, a novel Feature-Score Correlation-Aware Rehearsal is proposed to store and reuse data from previous tasks with limited memory size. Secondly, an Action General-Specific Graph is developed to learn and decouple the action-general and action-specific knowledge so that the task-consistent score-discriminative features can be better extracted across various tasks. Extensive experiments are conducted to evaluate the contributions of proposed components. The comparisons with the existing continual learning methods additionally verify the effectiveness and versatility of our approach.
@@ I Introduction
Action Quality Assessment (AQA) aims to assess the quality of action in videos. Rather than classifying the video samples by their action types [1, 2, 3], a common setting of AQA is to predict a subtle action quality score for a video action instance. Recently, AQA has attached more and more attention for its great potential value for sports skills analysis [4, 5, 6, 7], surgical maneuver training [8, 9, 10] and so on[11, 12].
Although remarkable progress has been achieved, a remaining challenge in AQA has yet to be explored. In practical application, the increasing demand drives the model to sequentially expand its applicability on different tasks (e.g., new technical actions or various daily actions). A more practical example is that, considering someday you may hope your AI cooking assistant can guide you on how to complete a French dish, but currently, it may only assess how well you cook Japanese cuisine. However, existing works on AQA assume that all the training data are visible for training at one time, ignoring such increasing demands. The already trained model may not work on the newly come tasks, resulting in misjudging or misguiding.
In this work, we consider a more complex but real scenario, Continual Learning (CL) in AQA, namely Continual-AQA. Like the other deep neural networks, AQA models also suffer from Catastrophic Forgetting[13, 14, 15]. As shown in Figure 1, when directly fine-tuning an already trained model from one AQA task to another, the performance significantly drops no matter adapting the model to a dissimilar task (e.g., Sync.3m and Skiing are quite not similar) or even a similar one (e.g., both Sync.3m and Single.10m are diving-like actions).
@@
To mitigate the forgetting, we start by analyzing the different learning targets between Conventional CL and Continual-AQA. As shown in Figure 2, we claim that to solve Continual-AQA, the key is to sequentially construct a task-consistent score-discriminative feature distribution across all seen tasks rather than learning disjoint feature clusters and decision boundaries for different classes or tasks in conventional CL [16, 17, 18]. A further explanation of this claim is that if an AQA model can map video actions of seen tasks to the feature space, in which the latent features express a strong correlation with the score labels regardless of the task or action types, it would be able to handle various tasks without forgetting. Moreover, considering different AQA tasks have general and specific assessing criteria (e.g., body balance and motion fluency are general criteria; landing distance and splash are specific criteria for Snowboarding and Diving, respectively), it is challenging to extract task-consistent score-discriminative features without explicitly modeling such general and specific knowledge. Therefore, to achieve Continual-AQA with less forgetting, we raise the two following problems: 1) How to fuse rather than isolate the previously learned feature distribution with the newly learned one? 2) How to explicitly model general and specific knowledge across tasks so that consistent score-discriminative features can be better extracted?
@@
To address the problems above, we propose a novel Feature-Score Correlation-Aware Rehearsal (FSCAR) to store and reuse data from previous tasks with limited memory size. First, to sample representative exemplars, an intuitive but effective Grouping Sampling strategy is adopted to sample exemplars that cover different score levels. Additionally, to address the problem that the model would prefer to learn data from the current task due to the imbalance of previous and current data [18, 19, 20], a Feature-Score co-Augmentation is proposed to augment the latent features and the scores of previous data. We introduce the concept of “augmentation helpers”, which provide information on the feature distribution of previous data so that the augmentation can be guided and the augmented feature distribution can correctly reflect the correlation between features and action quality scores. Moreover, an alignment loss based on difference modeling [21] is adopted to align the features across tasks and construct a task-consistent score-discriminative feature distribution.
Furthermore, we separately model the action-general and action-specific knowledge to better extract consistent score-discriminative features across different tasks). Specifically, we build an Action General-Specific Graph (AGSG) over the original joint relation graph proposed in [4]. AGSG decouples the original joint relation graph into two parts (i.e., Action General Graph and Action Specific Graph) to model action-general and action-specific knowledge, respectively. We observe that such a design can further mitigate forgetting.
In summary, the contributions of our works are as follows: 1) To our best knowledge, this is the first work to study Continual-AQA. 2) We propose a novel Feature-Score Correlation-Aware Rehearsal (FSCAR) to store and reuse data from previous tasks. Compared with existing rehearsal-based methods, FSCAR is more suitable for Continual-AQA with its feature-score correlation awareness. 3) We propose to learn and decouple action-general and action-specific knowledge with an Action General-Specific Graph module for learning different AQA tasks sequentially with less forgetting.
@@ II Related Works
Action Quality Assessment (AQA) aims to evaluate and quantify the overall performance or proficiency of human actions based on video or motion data analysis. Most works in AQA mainly focus on problem formulations [11, 25, 21, 6, 26], designing effective feature extractors [7, 12, 4, 27, 28, 29, 30], data insufficiency [31, 32], and so on. However, most existing methods only consider training the model at one time, ignoring the problem of training a model continually. In this paper, considering the characteristics of Continual-AQA, we attempt to find a feasible solution for this problem.
The closest to our work is [4] because our AGSG is built on the proposed JRG module. However, the original JRG is not designed for Continual Learning, and AGSG differs from it in graph calculation and model training. Additionally, a concurrent work [33] considers the transferable knowledge across various tasks and tries to model such knowledge explicitly, but it assumes that data from different tasks are all available during training, which contradicts the invisibility of previous data in Continual-AQA.
@@
To address catastrophic forgetting, Rehearsal-based methods propose to save and reuse several samples of older tasks. In the past few years, many methods [16, 18, 19, 20, 24] focused on selecting representative exemplars and balancing new and limited old data. However, compared with convention classification-based CL, Continual-AQA has a totally different learning target, which makes the existing methods [16, 18, 19, 20] unsuitable for Continual-AQA with their classification-based designs. In this work, a novel rehearsal-based method is proposed considering the distinct learning target explained in Figure 2. Recently, inspired by Complementary Learning, some works [34, 17] propose to model task-invariant and task-specific knowledge to address catastrophic forgetting. However, existing works either heavily rely on pre-training [34] or attempts to construct a disjoint representation for the private knowledge of each task [17]. In this work, we address Continual-AQA without any pre-training on other AQA tasks. Additionally, we design an Action General-Specific Graph module to extract general and specific knowledge to map the input video actions from different tasks to the features in a task-consistent action quality latent space.
@@ III Problem Formulation
We first review the general setting of AQA: given a video instance that contains one action (e.g., diving or skiing) to be assessed, the model is expected to predict the action quality score of the action. Following existing works [35, 4, 7, 36], we formulate AQA as a regression-based task.
Now considering a sequence of AQA tasks $D=\{D^{1},...,D^{T}\}$, the $t$-th task $D^{t}$ contains $N^{t}$ video instances $v^{t}$ and their action quality scores. Tasks with various action types (e.g., Diving and Snowboard) will arrive sequentially, and the model will be trained each time a new task arrives. When training the model on a new task $D^{t}$, data from previous tasks will not be available, or just a tiny amount (less than a predefined small number $M$) of them can be stored and reused. The goal of Continual-AQA is to train a unified model $F_{\theta}:V\to S$ parameterized by $\theta$, so that it can predict the action quality score $\hat{s}=F_{\theta}(v)\in S$ correctly when given an unseen test video action $v$ from any task has been seen before. In this work, we separate the model into a feature extractor $E$ and a lightweight score regressor $R_{s}$, which means $\hat{s}=F_{\theta}(v)=R_{s}(E(v))$.
Figure 3: A pipeline of our method for Continual-AQA. Our framework contains two main components: Feature-Score Correlation-Aware Rehearsal(FSCAR) in sub-image (b) and Feature Extractor with Action General-Specific Graph(AGSG) in sub-image (a). Notably, when a new task comes, the model is initialized by the previously learned weights, and a new Action-Specific Graph will be defined for the new task before training. An intuitive example of Feature-Score co-Augmentation is shown in sub-image (c), where a feature with a darker color has a higher quality score. Arrows in different colors indicate different data streams. $\{1,...,t\}$ is the index of tasks and here we denote task $t$ is the current task and task $t\text{-}1$ is the nearest previous task. $D^{t}$ denotes all the training data in current task, and $\psi=\{\psi^{1},...,\psi^{t-1}\}$ denotes the stored exemplars of previous tasks. Best viewed in color.
@@ IV-A Overview of Our Method
As illustrated in Section I, Continual-AQA has a different learning target from conventional Classification-based Continual Learning, that is, learning a task-consistent score-discriminative feature distribution across all seen tasks. Motivated by this, two main questions have been raised: 1) How to fuse rather than isolate the previously learned feature distribution with the newly learned one? 2) How to model the general and specific knowledge so that score-discriminative features can be better extracted across tasks? To address the two questions, we develop our method from the following two aspects.
@@ IV-B Feature-Score Correlation-Aware Rehearsal
In Continual-AQA, an ideal way to reuse insufficient previous data is to give an adequate description of the already learned feature distribution and fuse the feature of the old and new data into a task-consistent score-discriminative distribution so that the regressor can perform an accurate prediction on the scores regardless of the action types. To this end, we propose a Feature-Score Correlation-Aware Rehearsal, which mainly solves two problems: ”what to replay” and “how to replay” for Continual-AQA.
After training on task ${t\text{-}1}$, we adopt a Grouping Sampling (GS) strategy to sample representative exemplars of task $t\text{-}1$. Suppose we need to sample $m$ exemplars to store in the memory of size M and reuse them in the future training stages. GS first uniformly separates data into $m$ groups based on the score ranking. After that, one exemplar will be sampled from each group. The sampled exemplars of task $t\text{-}1$ are denoted as $\psi_{t-1}$. Although such a sampling strategy is intuitive, experiments in Section V-D show that it can already roughly describe the relationship between video action instances and quality scores and cover exemplars with different-level scores for previous data.
To reuse the stored data and solve the imbalance problem [18, 19, 20], when training on task $t$ ($t>1$), the stored exemplars of previous tasks $\{\psi^{1},...,\psi^{t-1}\}$ are fed into feature extractor $E^{t}$ and $E^{t-1}$. Suppose $f_{pre,j}^{t}$ and $f_{pre,j}^{t-1}$ denote the features of a previous instance extracted by $E^{t}$ and $E^{t-1}$. Then, a Feature-Score co-Augmentation(FS-Aug) is applied on $f_{pre,j}^{t-1}$ and its corresponding action quality score $s_{pre,j}$ to obtain the augmented feature $\tilde{f}^{t-1}_{pre,j}$ and score $\tilde{s}_{pre,j}$. Explicitly, given a feature to be augmented and its corresponding action quality score, FS-Aug first randomly selects K feature-score pairs from previous data as the augmentation helpers, which can be denoted as $H_{feat}=\{f_{hlp,1}^{t-1},...,f_{hlp,K}^{t-1}\}$ and $H_{score}=\{s_{hlp,1},...,s_{hlp,K}\}$. Then, the feature perturbation $\Delta f$ and the score perturbation $\Delta s$ are computed as:
@@
Finally, an auxiliary loss is adopted to align the features of previous and current data into a consistent distribution. Specifically, inspired by the idea of score difference modeling proposed in CoRe [21], an auxiliary difference regressor $R^{t}_{d}$ is utilized to predict the score difference between $\tilde{f}^{t-1}_{pre}$ and $f_{pre}^{t}$, and the difference loss is computed as:
@@
- Discussions. It has been emphasized in previous works [18, 19, 20] that the imbalance of previous and current data will cause forgetting. Hence, inspired by PASS [18], besides training $R^{t}_{s}$ with current data and stored previous data, we propose FS-Aug to augment previous data to address the imbalance issue. We have to claim that augmenting the corresponding scores together with the features is necessary because of the strong correlation between features and scores, which means a subtle perturbation in features would cause a perturbation in scores. Noting that PASS [18] augments feature by adding random Gaussian noise, it is impossible for PASS to perceive such a strong correlation. Differently, in our FS-Aug, augmentation helpers are introduced to provide information on the subtle correlation between latent features and scores so that the augmentation can be guided. Moreover, we adopt a difference modeling-based loss to maintain the score-discriminability of $\tilde{f}^{t-1}_{pre,j}$ and $f_{pre,j}^{t}$ given any $i$ so that the forgetting can be constrained. Ablation studies in Sec. V-D show that our approach outperforms the augmentation method proposed in PASS [18] and other possible strategies.

### v5

@@ Abstract
Action Quality Assessment (AQA) is a task that tries to answer how well an action is carried out. While remarkable progress has been achieved, existing works on AQA assume that all the training data are visible for training at one time, but do not enable continual learning on assessing new technical actions. In this work, we address such a Continual Learning problem in AQA (Continual-AQA), which urges a unified model to learn AQA tasks sequentially without forgetting. Our idea for modeling Continual-AQA is to sequentially learn a task-consistent score-discriminative feature distribution, in which the latent features express a strong correlation with the score labels regardless of the task or action types. From this perspective, we aim to mitigate the forgetting in Continual-AQA from two aspects. Firstly, to fuse the features of new and previous data into a score-discriminative distribution, a novel Feature-Score Correlation-Aware Rehearsal is proposed to store and reuse data from previous tasks with limited memory size. Secondly, an Action General-Specific Graph is developed to learn and decouple the action-general and action-specific knowledge so that the task-consistent score-discriminative features can be better extracted across various tasks. Extensive experiments are conducted to evaluate the contributions of proposed components. The comparisons with the existing continual learning methods additionally verify the effectiveness and versatility of our approach. Data and code are available at https://github.com/iSEE-Laboratory/Continual-AQA.
@@ I Introduction
Action Quality Assessment (AQA) aims to assess the quality of action in videos. Rather than classifying the video samples by their action types [1, 2, 3, 4, 5], a common setting of AQA is to predict a subtle action quality score for a video action instance. Recently, AQA has attached more and more attention for its great potential value for sports skills analysis [6, 7, 8, 9], surgical maneuver training [10, 11, 12] and so on[13, 14].
Although remarkable progress has been achieved, a remaining challenge in AQA has yet to be explored. In practical application, the increasing demand drives the model to sequentially expand its applicability on different tasks (e.g., new technical actions or various daily actions). A more practical example is that, considering someday you may hope your AI cooking assistant can guide you in preparing a French dish, but currently, it may only assess how well you cook Japanese cuisine. However, current AQA methods assume that all the training data is visible for training at one time, neglecting the necessity for continual learning to meet evolving demands. Consequently, the already trained model may not work on the newly come tasks, resulting in misjudging or misguiding.
In this work, we delve into a more complex but real scenario, Continual Learning (CL) in AQA, namely Continual-AQA. Similar to the other deep neural networks, AQA models also suffer from Catastrophic Forgetting[15, 16, 17]. As shown in Figure 1, when directly fine-tuning an already trained model from one AQA task to another, there is a noticeable decline in performance no matter adapting the model to a dissimilar task (e.g., Sync.3m and Skiing are quite not similar) or even a similar one (e.g., both Sync.3m and Single.10m are diving-like actions).
@@
To mitigate the forgetting, we begin with an analysis of the difference learning targets between Conventional CL and Continual-AQA. As shown in Figure 2, we claim that to solve Continual-AQA, the key is to sequentially construct a task-consistent score-discriminative feature distribution across all seen tasks rather than learning disjoint feature clusters and decision boundaries for distinct classes or tasks in conventional CL [18, 19, 20]. A further explanation of this claim is that if an AQA model can map action videos of seen tasks to the feature space, where the latent features express a strong correlation with the score labels regardless of the task or action types, it would be able to handle various tasks without forgetting. Moreover, considering different AQA tasks have general and specific assessing criteria (e.g., body balance and motion fluency are general criteria; landing distance and splash are specific criteria for Snowboarding and Diving, respectively), it is challenging to extract task-consistent score-discriminative features without explicitly modeling such general and specific knowledge. Therefore, to achieve Continual-AQA with less forgetting, we raise the two following problems: 1) How to fuse rather than isolate the feature distribution learned from current task with that learned from the previous tasks? 2) How to explicitly model general and specific knowledge across tasks so that consistent score-discriminative features can be better extracted?
@@
To address the aforementioned problems, we propose a novel Feature-Score Correlation-Aware Rehearsal (FSCAR) to efficiently utilize data from previous tasks within the constraint of limited memory size. First, to sample representative exemplars, an intuitive but effective Grouping Sampling strategy is adopted to sample exemplars that cover different score levels. Additionally, to address the imbalance between previous and current data [20, 21, 22], we propose Feature-Score co-Augmentation to augment the latent features and the scores of previous data. Specifically, we introduce the concept of “augmentation helpers”, which provide information on the feature distribution of previous data so that the augmentation can be guided and the augmented feature distribution can correctly reflect the correlation between features and action quality scores. Moreover, an alignment loss based on difference modeling [23] is adopted to align the features across tasks and construct a task-consistent score-discriminative feature distribution.
Furthermore, we separately model the action-general and action-specific knowledge to enhance the extraction of task-consistent score-discriminative features across different tasks). Specifically, we build an Action General-Specific Graph (AGSG) over the original joint relation graph proposed in [6]. AGSG decouples the original joint relation graph into two parts (i.e., Action General Graph and Action Specific Graph) to model action-general and action-specific knowledge, respectively. We observe that such a design can further mitigate forgetting.
In summary, the contributions of our work are as follows: 1) To our best knowledge, this is the first work to study Continual-AQA. 2) We propose a novel Feature-Score Correlation-Aware Rehearsal (FSCAR) to store and reuse data from previous tasks. FSCAR distinguishes itself from existing rehearsal methods by its awareness of feature-score correlations. 3) We propose to learn and decouple action-general and action-specific knowledge with an Action General-Specific Graph module for learning different AQA tasks sequentially with less forgetting.
@@ II Related Works
Action Quality Assessment (AQA) aims to evaluate and quantify the overall performance or proficiency of human actions based on video or motion data analysis. Most works in AQA mainly focus on problem formulations [13, 27, 23, 8, 28], designing effective feature extractors [9, 14, 6, 29, 30, 31, 32, 33], data insufficiency [34, 35], and so on. However, existing methods only consider training the model at one time, ignoring the problem of training a model continually. In this paper, considering the characteristics of Continual-AQA, we attempt to find a feasible solution for this problem.
The closest to our work is [6] because our AGSG is built on the proposed JRG module. However, the original JRG is not designed for Continual Learning, and AGSG differs from it in graph construction and model training. Additionally, a concurrent work [36] considers the transferable knowledge across various tasks and tries to model such knowledge explicitly, but it assumes that data from different tasks are all available during training, which contradicts the invisibility of previous data in Continual-AQA.
@@
To address catastrophic forgetting, Rehearsal-based methods propose to save and reuse several samples of older tasks. In the past few years, many methods [18, 20, 21, 22, 26] focused on selecting representative exemplars and balancing new and limited old data. However, compared with convention classification-based CL, Continual-AQA has a totally different learning target, which makes the existing methods [18, 20, 21, 22] unsuitable for Continual-AQA with their classification-based designs. In this work, a novel rehearsal-based method is proposed considering the distinct learning target explained in Figure 2. The most related works are iCaRL [18] and PASS [20]. The former one proposes to store limited instances of previous tasks, but the class-clustered based sampling strategy cannot sample the representative exemplars in Continual-AQA. The latter one proposes to randomly augment the stored features with a Gaussian noises, ignoring the subtle relation between features and scores inter- and intra-AQA tasks. The Feature-Score Correlation-Aware Rehearsal overcome the weakness of these methods by proposing a grouping sampling to sample representative exemplars and a feature-score co-augmentation to model the subtle relation across stored instances and thier corresponding scores so that the augmentation can be guided.
Recently, inspired by Complementary Learning [37, 38], some works [39, 19] propose to model task-invariant and task-specific knowledge to address catastrophic forgetting. However, existing works either heavily rely on pre-training [39] or attempts to construct a disjoint representation for the private knowledge of each task [19]. In this work, we address Continual-AQA without any pre-training on other AQA tasks. Additionally, we design an Action General-Specific Graph module to extract general and specific knowledge to map the input video actions from different tasks to the features in a task-consistent action quality latent space.
@@ III Problem Formulation
We first review the general setting of AQA: given an action video contains a single action (e.g., diving or skiing), the objective for the model is to accurately predict the quality score of the action. Following existing works [40, 6, 9, 32], we formulate AQA as a regression-based task.
Now considering a sequence of AQA tasks $D=\{D^{1},...,D^{T}\}$, each task $D^{t}$ comprises $N^{t}$ action videos $v^{t}$ with their corresponding action quality scores. These tasks, encompassing a variety of action types (e.g., Diving and Snowboarding), arrive sequentially. Upon the arrival of a new task, the model undergoes training on it. It is noteworthy that when training the model on a new task $D^{t}$, data from previous tasks will not be available, or just a tiny amount (less than a predefined small number $M$) of them can be stored and reused.
The goal of Continual-AQA is to train a unified model $F_{\theta}:V\to S$ parameterized by $\theta$, so that it can predict the action quality score $\hat{s}=F_{\theta}(v)\in S$ correctly for an unseen test action video $v$ from any task seen before. In this work, we separate the model into a feature extractor $E$ and a lightweight score regressor $R_{s}$, which means $\hat{s}=F_{\theta}(v)=R_{s}(E(v))$.
Figure 3: A pipeline of our method for Continual-AQA. Our framework contains two main components: Feature-Score Correlation-Aware Rehearsal(FSCAR) in sub-image (b) and Feature Extractor with Action General-Specific Graph(AGSG) in sub-image (a). Notably, when a new task comes, the model is initialized by the previously learned weights, and a new Action-Specific Graph will be defined for the new task before training. An intuitive example of Feature-Score co-Augmentation is shown in sub-image (c), where a feature with a darker color has a higher quality score. Arrows in different colors indicate different data streams. $\{1,...,t\}$ is the index of tasks and here we regard task $t$ as the current task and task $t\text{-}1$ as the nearest previous task. $D^{t}$ denotes all the training data in current task, and $\psi=\{\psi^{1},...,\psi^{t-1}\}$ denotes the stored exemplars of previous tasks. Best viewed in color.
@@ IV-A Overview of Our Method
As illustrated in Section I, Continual-AQA has a different learning target from conventional Classification-based Continual Learning, that is, learning a task-consistent score-discriminative feature distribution across all seen tasks. Motivated by this, two main questions have been raised: 1) How to fuse rather than isolate the feature distribution learned from previous tasks with that learned from the new task? 2) How to explicitly model general and specific knowledge across tasks so that task-consistent score-discriminative features can be better extracted? To address the two questions, we develop our method from the following two aspects.
@@ IV-B Feature-Score Correlation-Aware Rehearsal
Our goal is to ensure that the model can accurately represent the feature distribution of previous tasks using a limited stored data. This capability is crucial for integrating new data into this established distribution, resulting in a task-consistent, score-discriminative feature distribution. Such a distribution enables the regressor to reliably predict scores regardless of action types. To achieve this, we propose a Feature-Score Correlation-Aware Rehearsal, which is designed to tackle two problems, i.e., “what to replay” and “how to replay” for Continual-AQA.
Upon completing training on task ${t\text{-}1}$, we adopt a Grouping Sampling (GS) strategy to sample representative exemplars of task $t\text{-}1$. Suppose we need to sample $m$ exemplars to store in the memory of size M and reuse them in the future training stages. GS first uniformly separates data into $m$ groups based on the score ranking. Subsequently, one exemplar will be sampled from each group. The sampled exemplars of task $t\text{-}1$ are noted as $\psi_{t-1}$. Although such a sampling strategy is intuitive, experiments in Section V-D show that it can already roughly describe the relationship between video action instances and quality scores and cover exemplars with different-level scores for previous data.
To reuse the stored data and solve the imbalance problem [20, 21, 22], when training on task $t$ ($t>1$), the stored exemplars of previous tasks $\{\psi^{1},...,\psi^{t-1}\}$ are fed into feature extractor $E^{t}$ and $E^{t-1}$. Suppose $f_{pre,j}^{t}$ and $f_{pre,j}^{t-1}$ denote the features of a previous instance extracted by $E^{t}$ and $E^{t-1}$, where $j$ indicates the index of features of stored previous instances. Then, a Feature-Score co-Augmentation(FS-Aug) is applied on $f_{pre,j}^{t-1}$ and its corresponding action quality score $s_{pre,j}$ to obtain the augmented feature $\tilde{f}^{t-1}_{pre,j}$ and score $\tilde{s}_{pre,j}$. Specifically, given a feature to be augmented and its corresponding action quality score, FS-Aug first randomly selects K feature-score pairs from previous data as the augmentation helpers, which can be noted as $H_{feat}=\{f_{hlp,1}^{t-1},...,f_{hlp,K}^{t-1}\}$ and $H_{score}=\{s_{hlp,1},...,s_{hlp,K}\}$. Then, the feature perturbation $\Delta f$ and the score perturbation $\Delta s$ are computed as:
@@
After augmenting the features and scores of previous data, a remaining question is: how to utilize the augmented features and scores? To address this issue, our insight is, if the model can learn without forgetting, it should be easy to predict the score different between $\tilde{f}^{t-1}_{pre,j}$ and $f_{pre,j}^{t}$. Hence, an auxiliary loss is adopted to align the features of previous and current data into a consistent distribution. Specifically, inspired by the idea of score difference modeling proposed in CoRe [23], an auxiliary difference regressor $R^{t}_{d}$ is utilized to predict the score difference between $\tilde{f}^{t-1}_{pre}$ and $f_{pre}^{t}$, and the difference loss is computed as:
@@
- Discussions. It has been emphasized in previous works [20, 21, 22] that the imbalance of previous and current data will cause forgetting. Hence, inspired by PASS [20], besides training $R^{t}_{s}$ with current data and stored previous data, we propose FS-Aug to augment previous data to address the imbalance issue.
We have to claim that augmenting the corresponding scores together with the features is necessary because of the strong correlation between features and scores, which means a subtle perturbation in features would cause a perturbation in scores. Noting that PASS [20] augments feature by adding random Gaussian noise, it is impossible for PASS to perceive such a strong correlation. Differently, in our FS-Aug, augmentation helpers are introduced to provide information on the subtle correlation between latent features and scores so that the augmentation can be guided.
The deeper motivation of Equation 1-4 is also based on the strong correlation between features and scores. Imagine two features $f_{i}$ and $f_{j}$ in the latent space (corresponding action quality scores are noted as $s_{i}$ and $s_{j}$, and $s_{i}>s_{j}$), some feature point $f_{z}$ around the connection between $f_{i}$ and $f_{j}$, with high possibility, would corespond to a score $s_{z}$ which satisfies $s_{i}>s_{z}>s_{j}$. Hence, we can obtain the augmentation direction of $f^{t-1}_{pre,j}$ and score $s^{t-1}_{pre,j}$ with the help of other features (i.e., helpers) in the same latent space and its corresponding scores.
Based on FS-Aug, we adopt a difference modeling-based loss to maintain the score-discriminability of $\tilde{f}^{t-1}_{pre,j}$ and $f_{pre,j}^{t}$ given any $i$ so that the forgetting can be constrained. Ablation studies in Sec. V-D show that our approach outperforms the augmentation method proposed in PASS [20] and other possible strategies.